# Think Together and Work Better: Combining Humans’ and LLMs’ Think-Aloud Outcomes for Effective Text Evaluation

### v1

@@ Abstract.
This study introduces InteractEval, a framework that integrates human expertise and Large Language Models (LLMs) using the Think-Aloud (TA) method to generate attributes for checklist-based text evaluation. By combining human flexibility and reasoning with LLM consistency, InteractEval outperforms traditional non-LLM-based and LLM-based baselines across four distinct dimensions, consisting of Coherence, Fluency, Consistency, and Relevance. The experiment also investigates the effectiveness of the TA method, showing that it promotes divergent thinking in both humans and LLMs, leading to the generation of a wider range of relevant attributes and enhance text evaluation performance. Comparative analysis reveals that humans excel at identifying attributes related to internal quality (Coherence and Fluency), but LLMs perform better at those attributes related to external alignment (Consistency and Relevance). Consequently, leveraging both humans and LLMs together produces the best evaluation outcomes. In other words, this study emphasizes the necessity of effectively combining humans and LLMs in an automated checklist-based text evaluation framework. The code is available at https://github.com/BBeeChu/InteractEval.git.
@@
Figure 1. The overall process of prompting-based, checklist-based text evaluation models, and InteractEval (Ours)
@@ 1. Introduction
Nevertheless, concerns persist regarding the manually created key components in these checklists. As shown in Figure 1 (B), the checklists are generated by LLMs through sub-questions derived from key components manually defined by the authors based on a rubric. This process raises two main concerns. First, the reliance on human-defined key components introduces a strong dependency on human judgment. According to previous studies, humans and models each possess distinct strengths (Pantcheva, 2023; Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023). Thus, depending exclusively on human-defined key components could introduce bias into the evaluation process. Secondly, it raises the question of whether these components fully capture how the aspects intended to be evaluated in the rubrics’ rating dimensions manifest in the actual text. This approach follows a theory-based method (Plakans, 2013; Rodríguez, 2008), where a rating scale is designed based on established theories or pre-defined guidelines, such as rubrics. This method potentially overlooks the detailed features represented in example texts, focusing only on the rubric, and bring about a coarse-grained checklists (Brindley, 1998; Turner and Upshur, 2002).
@@
This study introduces InteractEval, addressing the limitations of relying solely on either humans or machines by combining their ideas to mitigate biases that arise from over-reliance on one source. We merge the ideas of both LLMs and human experts derived from the think-aloud (TA) process, known for fostering divergent thinking (Ericsson and Simon, 1980; Ericsson, 6; Nielsen, 1994; Lewis, 1982), with rubrics and sample texts. These combined TA results are used to create unbiased and fine-grained outcomes for text evaluation (see Figure 1). We expect this synergy to complement humans and LLMs as they hold different strengths, and reduce bias and improve checklist quality (Liu et al., 2023; Wang et al., 2024a). Human experts excel in flexible, high-level reasoning (Wu et al., 2023; Blattberg and Hoch, 1990; Pantcheva, 2023), while models offer consistent output, even when processing a large amounts of information (Wang et al., 2024a; Das et al., 2023; Blattberg and Hoch, 1990). During the TA process, we merge performance-driven approaches, which review individual samples, with the theory-based methods. This makes humans and LLMs specifically consider how the aspects intended to be rated in a dimension are manifested in sample texts (Fulcher et al., 2011; Knoch, 2009; Jeffrey, 2015).
@@
Our framework contains three main steps to combine the outcomes of TA process and create checklists based on them. First, after the TA by humans and LLMs, we collect thoughts from both humans and LLMs, referred to as text attributes, which are needed for rating specific dimensions of texts. Next, we prompt an LLM to extract key components from the merged ideas, to cluster the merged thoughts based on these components, and to generate questions related to each component. Lastly, we ask the LLM to create a checklist for rating a specific dimension of the given texts.
The research questions in this paper are:
@@
    Does our text evaluation framework, which uses checklists designed with integrated attributes from human experts and LLMs through the TA method, outperform baseline text evaluation methods in terms of correlation with human judgments?
@@
    Do the attributes generated through the TA method improve the evaluation performance of generated texts, based on the checklists constructed from those attributes?
@@
    How effective is the combination of attributes from humans-TA and LLMs-TA for text evaluation using the checklists constructed from the TA-generated attributes?
@@ 2.1. Text Evaluation with Large Language Models
LLM-based text evaluation has emerged to address the shortcomings of traditional text evaluation methods, such as BLUE (Papineni et al., 2002) or ROUGE-L (Lin, 2004), which highly depend on reference text (Fabbri et al., 2021; Gu et al., 2021; Guan et al., 2021). Research on text evaluation using LLMs (Liu et al., 2023; Fu et al., 2023; Kocmi and Federmann, 2023; Chan et al., 2023; Kim et al., 2024) has been actively progressing, starting with GPTScore (Fu et al., 2023) which score texts based on its token generation probability and G-Eval (Liu et al., 2023) which use the Chain-of-Thought (CoT) (Wei et al., 2022) technique to evaluate texts. Studies utilizing multi-agent collaboration (Chan et al., 2023; Kim et al., 2024) are particularly active, aiming to overcome the single-thread limitations (Wang et al., 2024b) of single-LLM approaches, which hinder scalability and speed when dealing with complex tasks like text evaluation, which requires taking multiple aspects of text quality into consideration. Despite these advantages, those methods are resource-intensive, incurring high costs in time and computation. They experience increased time to first token (TTFT) (Wang et al., 2024b), and still fall short in their robustness on text evaluation. CheckEval (Lee et al., 2024b) introduced a binary evaluation method using checklists to assess texts, based on the observation that LLMs excel at fine-grained tasks (Liu et al., 2022; Min et al., 2023) and are particularly effective for yes/no questions. However, the heavy reliance on humans (specifically the authors) in designing key components for checklist creation raises concerns about the reliability of the evaluation criteria (Lee et al., 2024b). Moreover, it still suffers from being coarse-grained by designing the key components solely based on the broad rubric definition of each dimension.
@@ 2.2. Combination of Human Intelligence and Artificial Intelligence
Numerous studies have demonstrated that AI decision-making, when combined with human interventions, outperforms decisions in isolation (Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023; Revilla et al., 2023; Cohen et al., 2023). Historically, research has shown that human involvement improves the accuracy of statistical models in forecasting (Blattberg and Hoch, 1990; Clemen, 1989). Today, various studies continue to explore the effectiveness of human intervention in AI across different tasks. For example, in the retail industry, human involvement can significantly enhance an AI model’s predictions, especially when dealing with long time horizons and low uncertainty. Additionally, collaboration between humans and large language models (LLMs) improves the accuracy of data annotation (Wang et al., 2024a). LLM performance, notably, is highly dependent on specific tasks, datasets, and labels (Zhu et al., 2023; Ziems et al., 2024), and they often retain biases and errors in their outputs (Liu et al., 2023; Wang et al., 2024a), making human oversight crucial. Moreover, human-centered AI technologies remain essential for tasks like fact-checking (Das et al., 2023), highlighting the growing importance of human-AI interaction, particularly in high-level tasks like text evaluation within LLMs (Lee et al., 2024b).
Human experts and models each have unique strengths and weaknesses, making them complementary when used together. Humans are more flexible but can become fatigued or bored with large tasks, while models are more consistent and rigid (Blattberg and Hoch, 1990). Given the extensive prior knowledge (e.g. universal grammar (Montague et al., 1970; Chomsky et al., 2023)) humans possess, many studies in the human-in-the-loop (HITL) field aim to enhance the consistent performance of neural models by incorporating human expertise into data processing, model training, inference, and system development (Wu et al., 2022; Xin et al., 2018). Human elements, such as empirical knowledge and practical skills, help models learn and perform specific tasks, especially when dealing with sparse data (Wu et al., 2022; Kumar et al., 2019). Considering the biases in LLM performance (Liu et al., 2023; Wang et al., 2024a), human intervention can help mitigate these issues. Since most LLMs are trained with human feedback (Achiam et al., 2023; Touvron et al., 2023; Anthropic, 2024), and tasks like text generation and summarization can be improved with HITL (Ziegler et al., 2019; Stiennon et al., 2020), involving humans in checklist construction for text evaluation could enhance the quality of LLMs’ text ratings.
@@ 2.3. Think Aloud
The think aloud (TA) method, where participants verbalize their thoughts while performing experimental tasks, is a well-established technique in cognitive psychology and usability studies. Introduced by Ericsson and Simon (Ericsson and Simon, 1980), this method has been pivotal in understanding cognitive processes and identifying usability issues by capturing real-time user feedback (Nielsen, 1994; Lewis, 1982; Fan et al., 2019). There are two main types in the TA method: concurrent TA (CTA) and retrospective TA (RTA) (Ericsson, 6). In CTA, users verbalize their thoughts while performing a task, whereas in RTA, they recall and describe their thoughts and feelings after completing the task. Although both methods have their strengths, they also have limitations. To address them, a hybrid (HB) method has been developed that combines elements of both approaches (Alhadreti and Mayhew, 2018).
The TA method has been widely applied beyond usability studies, serving as a tool across various fields to develop new methodologies by capturing diverse ideas (Ericsson and Simon, 1980; Nielsen, 1994). For example, it was used to construct a rating scale for English writing assessment by having in-service teachers and experts think aloud their thoughts on relevant text attributes while reviewing sample essays and scoring dimensions (Wu, 2022). These insights were then consolidated into a detailed rubric, providing a structured approach to rubric creation. Building on (Wu, 2022), this study seeks to create a fine-grained rubric-based checklist for evaluating generated-texts. Using the TA method, we involve both humans and LLMs to identify specific text attributes for the rating dimensions-Coherence, Fluency, Consistency, and Relevance-ultimately refining and improving LLM-based evaluation.
@@ 3. Method
Figure 2. InteractEval Framework for Text Evaluation: (A) Think Aloud: Human experts and LLMs verbalize their thoughts to generate text attribute insights using sample texts and evaluation rubrics. (B) Checklist Construction: Insights are combined and categorized into key components, leading to the generation and validation of checklist questions. (C) Checklist-based Evaluation: The checklists are answered by an evaluator LLM to evaluate the summaries, with results aggregated into a final score.
@@ 3.1. Overview
InteractEval combines the TA outcomes from humans and LLMs to build checklists for reliable text evaluation. Inspired by the competitive and robust performance of CheckEval (Lee et al., 2024b) in text evaluation, this study adopts a checklist-based evaluation approach. InteractEval consists of three distinct stages, as illustrated in Figure 2. Initially, humans and LLMs participate in the TA process by independently brainstorming the text attributes to be considered when rating texts. These attributes are then merged and used to create checklists consisting of multiple boolean-style questions, requiring binary responses (Yes/No). The evaluator LLM reviews the text summaries based on the checklist and provides a response to each question. Finally, the ’Yes’ responses are tallied to generate a final score, which is then compared with the human-assigned ground truth scores. A key distinction from CheckEval is that InteractEval generates the components using the attributes generated by both human and LLM, reducing bias and including more fine-grained criteria (Pantcheva, 2023; Blattberg and Hoch, 1990; Brindley, 1998; Turner and Upshur, 2002).
@@ 3.2. Dimensions for Evaluation
The study uses SummEval (Fabbri et al., 2021) dataset, which includes a collection of news article summaries, for a case study. The dataset comprises source news articles on various topics, reference summaries, machine-generated summaries from different models, and human-evaluated quality scores (from 1.0 to 5.0) across four dimensions: Coherence, Fluency, Consistency, and Relevance. According to previous studies (Zhu and Bhat, 2020; Wang et al., 2024c), generated texts possess two key features: internal quality and external alignment. Internal quality refers to grammatical accuracy, readability, and structure, focusing on the intrinsic characteristics of the text. In contrast, external alignment evaluates adherence to prompts and factual accuracy, assessing how well the summary reflects and aligns with the source text. Based on the rubric from the original paper (Fabbri et al., 2021), we categorize Coherence and Fluency under internal quality, and Consistency and Relevance under external alignment. The detailed definition of each dimension is as follows:
  • Internal Quality
    • Coherence: The collective quality of all sentences. We align this dimension with the DUC quality question of structure and coherence whereby ”the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to a coherent body of information about a topic.
    • Fluency: The quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.
  • External Alignment
3.3. Attributes Collection through Think Aloud
@@
Figure 3. Human experts and LLMs are provided with dimension rubrics, sample texts for each dimension, and their corresponding scores. They then review the materials simultaneously, offering their thoughts regarding text attributes that should be considered when rating the dimension based on their knowledge.
In this stage, we collected the text attributes for constructing checklists of each dimension. For this, we utilized the TA method, which enhance divergent thinking (Ericsson, 6; Lewis, 1982). During the TA, participants verbalized their thoughts while conducting experimental tasks (Ericsson and Simon, 1980). Likewise, four human participants and four different LLM participants articulated the text attributes needed for rating the text quality of each dimension in this study. To minimize biases arising from individual members of each participant group, whether humans or LLMs, we included four members in each group. They did this while reading the definitions of each dimension along with four sample summaries per dimension as illustrated in Figure 3. As a result, the TA participants generated fine-grained attributes by considering the detailed features of each dimension as manifested in the example summaries. Between the CTA and the RTA (Ericsson, 6), which are two main methods of the TA, the current framework used the CTA, as it is straightforward to implement (Fan et al., 2019), avoids biases from post-task retrospection (Van den Haak et al., 2004; McDonald et al., 2012), and has been shown to have minimal impact on participants’ behavior (Hertzum et al., 2009).
3.3.1. TA by human experts
The human experts include two journalism and two linguistics specialists, chosen based on the nature of the news article summary dataset. We assumed that journalism experts would focus on aspects related to news content, while linguistics experts would concentrate on the language features of the texts. Their details are provided in Table 1. The experts conducted the process in real-time via online video calls, addressing each dimension—Coherence, Fluency, Consistency, and Relevance—separately. Specifically, we collected participants’ thoughts on each dimension through the following process. First, participants were provided with four news articles and summaries, actual scores corresponding to different score ranges, and the definitions of each dimension. Then, participants reviewed the news summaries corresponding to each score range and verbalized their reasoning for why they believe the news articles received the given scores. Finally, the participants’ verbalizations were recorded, transcribed, and the identified text attributes were compiled.

### v2

@@ Abstract.
This study introduces InteractEval, a framework that integrates the outcomes of Think-Aloud (TA) conducted by humans and LLMs to generate attributes for checklist-based text evaluation. By combining humans’ flexibility and high-level reasoning with LLMs’ consistency and extensive knowledge, InteractEval outperforms text evaluation baselines on a text summarization benchmark (SummEval) and an essay scoring benchmark (ELLIPSE). Furthermore, an in-depth analysis shows that it promotes divergent thinking in both humans and LLMs, leading to the generation of a wider range of relevant attributes and enhancement of text evaluation performance. A subsequent comparative analysis reveals that humans excel at identifying attributes related to internal quality (Coherence and Fluency), but LLMs perform better at those attributes related to external alignment (Consistency and Relevance). Consequently, leveraging both humans and LLMs together produces the best evaluation outcomes, highlighting the necessity of effectively combining humans and LLMs in an automated checklist-based text evaluation.
@@
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713181isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Human computer interaction (HCI)
Figure 1. Comparative Overview of Prompt-based, Checklist-based, and the Proposed InteractEval Text Evaluation Methods: (A) Prompt-based evaluation generates task-specific prompts from a rubric for prediction scores. (B) Checklist-based evaluation uses human-created questions from the rubric to form prompts. (C) InteractEval combines human and LLM-generated attributes using Think Aloud methods to create questions and produce final prediction scores.
@@ 1. Introduction
Nevertheless, concerns persist regarding the manually created key components in these checklists. As shown in Figure 1 (B), the checklists are generated by LLMs through sub-questions derived from key components manually defined by the authors based on a rubric. This process raises two main concerns. First, the reliance on human-defined key components introduces a strong dependency on human judgment. According to prior studies, humans and models each possess distinct strengths and weaknesses (Pantcheva, 2023; Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023). Thus, depending exclusively on human-defined key components may inherit the limitations of human abilities in the evaluation process. Secondly, it raises the question of whether these components fully capture how the aspects intended to be evaluated in the rubrics’ rating dimensions manifest in the actual text. This approach follows a theory-based method (Plakans, 2013; Rodríguez, 2008), where a rating scale is designed based on established theories or pre-defined guidelines, such as rubrics. This method potentially overlooks the detailed features represented in example texts, focusing only on the rubric, and bring about a coarse-grained checklists (Brindley, 1998; Turner and Upshur, 2002).
@@
Our study introduces InteractEval, addressing the limitations of relying solely on either humans or machines by combining their ideas to mitigate shortcomings that arise from over-reliance on one source. We integrate the ideas of LLMs and humans derived from think-aloud (TA) processes, which have been recognized for fostering divergent thinking (Ericsson and Simon, 1980; Ericsson, 6; Nielsen, 1994; Lewis, 1982). During the TA procedure, we merge performance-driven approaches, which review individual samples, with the theory-based methods, inducing humans and LLMs to specifically consider how the aspects intended to be rated in a dimension are reflected in sample texts (Fulcher et al., 2011; Knoch, 2009; Jeffrey, 2015). These combined TA results are then used to create unbiased and fine-grained checklists for text evaluation as illustrated in Figure 1 (C).
We expect the human-LLM collaboration to complement their distinct advantages via the combination (Liu et al., 2023; Wang et al., 2024a). Specifically, humans often struggle with inconsistent performance when handling large tasks, which can significantly hinder their performance (Li et al., 2020). In text evaluation, humans often find it challenging to identify relevance between two documents due to the need for consistent reading of large volumes of text. On the other hand, automated models heavily rely on extensive datasets (Wang et al., 2024a; Das et al., 2023), which restrict them to the knowledge and patterns explicitly present in the training data, often making it difficult for them to fully capture the complex semantic details of a given text. For example, models may encounter difficulties to identify coherent and hierarchical structure in a given text (Pantcheva, 2023), highlighting the models’ vulnerability to the quality and diversity of the data they are trained on. Therefore, integrating the consistent performance of models with humans’ flexibility, prior knowledge, and high-level reasoning capability (e.g., critical thinking, creativity, imagination, diverse perspective-taking, situational awareness, and intuitive thinking) can enhance the reliability of text evaluation by producing results more closely aligned with human judgement (Blattberg and Hoch, 1990; Das et al., 2023; Hong et al., 2022; Wu et al., 2023; Pantcheva, 2023).
@@
Our framework contains three main steps to combine the outcomes of TA and create checklists based on them. First, after the TA produres conducted by humans and LLMs, we collect the results from both of them, referred to as text attributes, which are needed for rating specific dimensions of texts. Next, we prompt an LLM to extract key components from the merged ideas, to cluster the merged attributes based on these components, and to generate questions related to each component. Lastly, we employ the LLM to create a checklist for rating a specific dimension of the given texts.
This study examines whether combining TA results from both humans and LLMs improves text evaluation performance. We conduct extensive experiments to test whether the TA method we use enables humans and LLMs to generate diverse attributes that contribute to building a refined checklist and enhancing evaluation quality. Additionally, we explore how the checklist, based on these combined attributes, further improves text evaluation. In sum, our study examines three key research questions, as outlined below:
@@
    Does our text evaluation framework, which uses checklists designed with integrated attributes from human experts and LLMs through TA, outperform baseline text evaluation methods in terms of correlation with ground-truth scores?
@@
    Do the attributes generated through TA improve the evaluation performance of generated texts, based on the checklists constructed from those attributes?
@@
    How effective is the combination of attributes from humans’ TA and LLMs’ TA for text evaluation using the checklists constructed from the TA-generated attributes?
@@ 2.1. Text Evaluation with Large Language Models
LLM-based text evaluation has emerged to address the shortcomings of traditional text evaluation methods, such as BLUE (Papineni et al., 2002) or ROUGE-L (Lin, 2004), which highly depend on reference text (Fabbri et al., 2021; Gu et al., 2021; Guan et al., 2021). Research on text evaluation using LLMs (Liu et al., 2023; Fu et al., 2023; Kocmi and Federmann, 2023; Chan et al., 2023; Kim et al., 2024) has been actively progressing, starting with GPTScore (Fu et al., 2023), which scores texts based on its token generation probability, and G-Eval (Liu et al., 2023), which uses the Chain-of-Thought (CoT) (Wei et al., 2022) technique to evaluate texts. Studies utilizing multi-agent collaboration (Chan et al., 2023; Kim et al., 2024) are particularly active, with the aim of overcoming the single-thread limitations (Wang et al., 2024b) of single-LLM approaches. The single-LLM methods hinder scalability and speed when dealing with complex tasks like text evaluation, which requires taking multiple aspects of text quality into consideration. Despite these advantages, those multi-agent based methods are resource-intensive, incurring high costs in time and computation. They experience increased time to first token (TTFT) (Wang et al., 2024b), and still fall short in their robustness on text evaluation. CheckEval (Lee et al., 2024b), which are particularly effective for yes/no questions, introduced a binary evaluation method using checklists to assess texts, based on the observation that LLMs excel at fine-grained tasks (Liu et al., 2022; Min et al., 2023). However, its heavy reliance on humans (specifically the authors) in designing key components for checklist creation raises concerns about the reliability of the evaluation criteria (Lee et al., 2024b). Moreover, it still suffers from being coarse-grained by designing the key components solely based on the broad rubric definition of each dimension.
@@ 2.2. Combination of Human Intelligence and Artificial Intelligence
Human experts and models each have unique strengths and weaknesses, making them complementary when used together. Humans are more flexible but can become fatigued or bored with large tasks, while models are more consistent and rigid (Blattberg and Hoch, 1990). Given the substantial prior knowledge (e.g. universal grammar (Montague et al., 1970; Chomsky et al., 2023)) humans possess, many studies in the human-in-the-loop (HITL) field aim to enhance the consistent performance of neural models by incorporating human expertise into data processing, model training, inference, and system development (Wu et al., 2022; Xin et al., 2018). Human elements, such as empirical knowledge and practical skills, help models learn and perform specific tasks, especially when dealing with sparse data (Wu et al., 2022; Kumar et al., 2019). On the other hand, models are capable of managing demanding workloads efficiently, although they heavily rely on datasets. Specifically, LLMs can analyze vast text corpora of academic papers to find a new research insight (Oppenlaender and Hämäläinen, 2023). Additionally, incorporating LLMs into the research cycle accelerates the overall process and enhances efficiency. However, this approach may face challenges such as hallucinated outputs, misalignment with established referent knowledge, and self-contradictions (Elagroudy et al., 2024). LLM performance, notably, is highly dependent on specific tasks, datasets, and labels (Zhu et al., 2023; Ziems et al., 2024), and they often retain biases and errors in their outputs (Liu et al., 2023; Wang et al., 2024a), making human oversight crucial.
Numerous studies have demonstrated that combining human and AI for decision-making outperforms decisions in isolation (Blattberg and Hoch, 1990; Wang et al., 2024a; Das et al., 2023; Revilla et al., 2023; Cohen et al., 2023). Various studies have explored the effectiveness of collaboration between humans and AI models across various tasks, including text generation and non-text object generation, but its potential in text evaluation remains underexamined. For example, synergistic collaboration between humans and LLMs, leveraging human cognitive strengths, produces higher-quality outputs in co-writing (Dhillon et al., 2024). Additionally, a machine-in-the-loop approach can enhance the workflow of human analyst’s text analysis by utilizing human knowledge as a foundation (Hong et al., 2022). Highly refined prompt templates, developed through the aggregation of diverse human expertise, can be employed to make LLMs precisely audit their own outputs (Rastogi et al., 2023), and collaboration between humans and LLMs supports the creative processes of individual designers or design teams for generating products (Zhou et al., 2024; Han et al., 2024). Furthermore, humans and LLMs can cooperate to improve the accuracy of data annotation (Wang et al., 2024a). Because most LLMs are trained with human feedback (Achiam et al., 2023; Touvron et al., 2023; Anthropic, 2024), and tasks like text generation and summarization can be improved with HITL (Ziegler et al., 2019; Stiennon et al., 2020), involving humans in checklist construction should enhance the quality of LLMs’ text evaluation.
@@ 2.3. Think Aloud
The think aloud (TA) method, where participants verbalize their thoughts while performing experimental tasks, is a well-established technique in cognitive psychology and usability studies. Introduced by Ericsson and Simon (Ericsson and Simon, 1980), this method has been pivotal in understanding cognitive processes and identifying usability issues by capturing real-time user feedback (Nielsen, 1994; Lewis, 1982; Fan et al., 2019). There are two main types in the TA method: concurrent TA (CTA) and retrospective TA (RTA) (Ericsson, 6). In CTA, users verbalize their thoughts while performing a task, whereas in RTA, they recall and describe their thoughts and feelings after completing the task. Although both methods have their own strengths, they also have limitations (Alhadreti and Mayhew, 2018).
The TA method has been widely applied beyond usability studies, serving as a tool across various fields to develop new methodologies by capturing diverse ideas (Ericsson and Simon, 1980; Nielsen, 1994). For example, it was used to construct a rating scale for English writing assessment by having in-service teachers and experts think aloud their thoughts on relevant text attributes while reviewing sample essays and scoring dimensions (Wu, 2022). These insights were then consolidated into a detailed rubric, providing a structured approach to rubric creation. Building upon the prior work, our study seeks to create a fine-grained checklist for evaluating texts. Using the TA method, we involve both humans and LLMs to identify specific text attributes for the rating dimensions - Coherence, Fluency, Consistency, and Relevance - ultimately refining and improving LLM-based evaluations.
@@ 3. Method
Figure 2. InteractEval Framework for Text Evaluation: (A) Think Aloud: Human experts verbalize their thoughts and LLMs articulate their knowledge to generate text attributes insights using sample texts and evaluation rubrics. (B) Checklist Construction: Insights are combined and categorized into key components, leading to the generation and validation of checklist questions. (C) Checklist-based Evaluation: The checklists are answered by an evaluator LLM to evaluate the summaries, with results aggregated into a final score, which is then checked against a ground-truth score.
@@ 3.1. Overview
InteractEval combines the TA outcomes from humans and LLMs to build checklists for reliable text evaluation. Inspired by the competitive and robust performance of CheckEval (Lee et al., 2024b) in text evaluation, this study adopts a checklist-based evaluation approach. InteractEval consists of three distinct stages, as illustrated in Figure 2. Initially, humans participate in the TA process, while LLMs are prompted in a manner identical to TA, as shown in Figure 3, independently producing the text attributes to be considered when rating texts. These attributes are then merged and used to create checklists consisting of multiple boolean-style questions, requiring binary responses (Yes/No). To rigorously and efficiently assess the impact of human-AI collaboration, we limit its scope to the TA stage, specifically evaluating how collaborative attributes generated during TA contribute to the checklist-based evaluation framework. The evaluator LLM reviews the text summaries based on the checklist and provides a response to each question. Finally, the ”Yes” responses are tallied to generate a final score, which is then compared with the human-assigned ground truth scores. A key distinction from CheckEval is that InteractEval generates the components using the attributes generated by both human and LLM, reducing shortcomings associated with each source and producing more fine-grained criteria (Pantcheva, 2023; Blattberg and Hoch, 1990; Brindley, 1998; Turner and Upshur, 2002).
@@ 3.2. Dimensions for Evaluation
Our study uses SummEval (Fabbri et al., 2021) dataset, which includes a collection of news article summaries, for the main case study. The dataset comprises source news articles on various topics, reference summaries, machine-generated summaries from different models, and human-evaluated quality scores (from 1.0 to 5.0) across four dimensions: Coherence, Fluency, Consistency, and Relevance, which are used as ground-truth scores in our study. According to prior studies (Zhu and Bhat, 2020; Wang et al., 2024c), generated texts commonly possess two key features: internal quality and external alignment. Internal quality refers to grammatical accuracy, readability, and structure, focusing on the intrinsic characteristics of the text. In contrast, external alignment evaluates adherence to prompts and factual accuracy, assessing how well the summary reflects and aligns with the source text. Based on the rubric from the original paper (Fabbri et al., 2021), we categorize Coherence and Fluency under internal quality, and Consistency and Relevance under external alignment. The detailed definition of each dimension is as follows (Fabbri et al., 2021):
@@
Figure 3. Text Attribute Generation Process: Human experts and LLMs are provided with a system message, a target dimension rubric, sample texts for the dimension, and their corresponding scores. They then review the materials, offering their ideas regarding text attributes that should be considered when rating the dimension based on their knowledge. Technically speaking, humans verbalize text attributes aloud through the process of thinking, whereas large language models (LLMs) generate text attributes after processing input prompt.